CLIMM(7) CLIMM(7) NAME climm−interactivecommands DESCRIPTION Thismanualpagedescribesthecommandsthatcanbeusedwithinclimm.Thispageholdsthe sameinformationastheonlinehelp,butmorepreciseandlonger.Notations:squarebrackets([]) denoteoptionalarguments,whileanglebrackets(<>)denotearguments(inoppositetokeywords).Specifically,<uin>denotesanarbitraryUIN,while<contacts>denotesacommaseparatedlistofcontacts,thatis,ofUINsandnicknames.Ifitisthelastargument,thecontactsmay alsobeseparatedbyspaces. COMMANDSCONCERNINGMESSAGESENDING msg<contacts>[<message>] Sendamessagetoallcontacts.Withoutmessage,orwithmessageendinginabackslash (\),multi−linemessagemodeisenabled.Tosendthemulti−linemessage,putaperiod (.)onalineofitsown.Tocancelthemessage,putahashsign(#)onalineofitsown. Outgoingandincomingmessagesaremarkeddifferentlydependingofthetransportused. <<<<<<and>>>>>>markmessagessentviaadirectconnection.<<%<<and>>%>>mark messagessentviaadirectconnectionthatwereencrypted(requiresGnuTLSorOpenSSL enabledatcompiletime).<<<<and>>>>markicq8(alsocalledtype−2)messages. Lastly,<<<and>>>markicq5messages,whichcanbesocalledtype−1,type−4or offlinemessage.Thesekindofmessagesarenotacknowledged. Amessageisfirsttriedtobesentviaapre−existingdirectconnection;ifnoneisestablished,oneisinitiatedinthebackgroundforfurthermessages.Ifnodirectconnectionis open,orifthemessageisnotacknowledgedaftersometime,theconnectionwillbefailed andthenexttransporttried:sendingastype−2message.Thisstepwillbeskippedifthe peerdoesnotsettherequiredcapabilitiestosignalitsabilitytoreceivethem.Ifthisis skipped,timesoutorreturnsanerror,e.g.thepeerhasgoneofflineinthetimebetween, themessageissentasanordinarytype−4message. Ifthepeersignalsitsabilitytoacceptthem,messagesaresentUTF−8encodedand appropriatelytagged.Otherwise,themessageissentencodedinthepeer´ sconfigured encoding,or,ifunset,thedefaultassumedremoteencoding.type−1,type−4andoffline messagescannotbetaggedwiththeirencoding,sosuccessfultransmissionof non−USASCII−charactersreliesoncorrectconfigurationandheuristics. a[<message>] Sendamessagetothelastpersonyousentamessageto.Auto-expandingaliasfor msg%a. r[<message>] Replytolastmessagereceived.Auto-expandingaliasformsg%r. url<contacts><url><message> Sendamessageregardinganurltoallcontacts. sms[<nick>][<cell>]<message> SendanSMSmessagemessagetocellphonenumbercell,whichlookslike+<country code><number>,ortothecellphonenumberofnick,whichwillbesettocellifunset. Youmaynotspecifycellifnickalreadyhasacellphonenumbergiven.Invalidcellphone numberswillberemovedfromthenick´ smetadata. chat<contacts>[<message>] Sendsamessagemessagetocontactscontactsjustasthemsgcommand.Afterthemessageissent(orcancelled),multi-linemodeisenteredandmessagessentuntilanempty messageiscancelled.Cancellinganon-emptymessagewillonlycancelthemessageitself, andmoremessagesinmulti-linemodecanbesent. climm 1 CLIMM(7) CLIMM(7) getauto[auto|away|na|dnd|occ|ffc][<contacts>] Requestsautomessagefromcontactsforgivenstatus,whereawayisforaway,nafornot available,dndfordonotdisturb,occforoccupiedandffcforfreeforchat.Ifautoor nothingisgiven,theautomessageforthecontactsisfetchedfortheirrespectivestatus. Contactsnotinanyofthosestatuswillbeskipped.Newfor0.4.10. auth[req|grant|deny|add]<contacts> Grantordenyallcontactstoaddyoutotheirlist,requestfromallcontactstoallowyou toaddthemtoyourcontactlist,ortellcontactsthatyouaddedthemtoyourcontact list. resend<contacts> Resendthelastmessagetomorecontacts. last[<contacts>] Showthelastmessagereceivedfromcontact,orfromeveryonewhoalreadysentamessage. h<contact>[<last>[<count>]] history<contact>[<last>[<count>]] Showmessagesfrom<contact>fromthelogfiles.Startatthelastlatestmessage,orthe -lastmessageiflastisnegative,defaultingtothe20iethlastmessage,anddisplaycount messages,defaultingto20ifnotgiven. historyd<contact|*><date>[<count>] Showmessagesfrom<contact>orallcontacts(for*)fromthelogfiles.Startatdate date,whichmustbegiveninISO8601format,eitherYYYY−MM−DDor YYYY−MM−DDTHH:MM:SS,anddisplaycountmessages,defaultingto20ifnotgiven. find<contact><pattern> Searchthepatternpatterncase−insensitiveinthelogfileforcontact. finds<contact><pattern> Searchthepatternpatterncase−sensitiveinthelogfileforcontact. tabsShowsofalistofusersyoucantabthrough. COMMANDSCONCERNINGYOURSTATUS loginAbuilt−inshortcutforconnlogin.Opensthefirstserverconnection. online[for<contacts>[<message>]] Changestatustoonline,orchangestatustoonlineonlyforthelistedcontactsand optionallyspecifythestatusmessageforthem. away[for<contacts>][<message>] Changestatustoawayandoptionallysetautoresponsetemporarelyforthisstatusto message,orchangestatusonlyforgivencontactsandoptionallyspecifythestatusmessageforthem. na[for<contacts>][<message>] Changestatustona(notavailable)andoptionallysetautoresponsetemporarelyforthis statustomessage,orchangestatusonlyforgivencontactsandoptionallyspecifythestatusmessageforthem. occ[for<contacts>][<message>] Changestatustoocc(occupied)andoptionallysetautoresponsetemporarelyforthisstatustomessage,orchangestatusonlyforgivencontactsandoptionallyspecifythestatus messageforthem. dnd[for<contacts>][<message>] Changestatustodnd(donotdisturb)andoptionallysetautoresponsetemporarelyfor thisstatustomessage,orchangestatusonlyforgivencontactsandoptionallyspecifythe statusmessageforthem. climm 2 CLIMM(7) CLIMM(7) ffc[for<contacts>][<message>] Changestatustoffc(freeforchat)andoptionallysetautoresponsetemporarelyforthis statustomessage,orchangestatusonlyforgivencontactsandoptionallyspecifythestatusmessageforthem. inv[for<contacts>][<message>] Changestatustoinv(invisible)andoptionallysetautoresponsetemporarelyforthisstatustomessage,orchangestatusonlyforgivencontactsandoptionallyspecifythestatus messageforthem. change[<number>[for<contacts>][<message>]] Withoutanumberitlistssomeavailablemodes.Otherwise,changestatustonumberand optionallysetautoresponsetemporarelyforthisstatustomessage,orchangestatusonly forgivencontactsandoptionallyspecifythestatusmessageforthem. COMMANDSCONCERNINGFINDINGANDSEEINGOTHERUSERS f<contacts> finger<contacts> Showallwhitepageinformationforcontacts<contacts>. ss<contacts> Showallsavedwhitepageinformationforcontacts<contacts>. iListallthepeopleonyourignorelist. s[<contacts>] Showyourcurrentstatus,orofallgivencontactsindetail,includingallaliases. e,ee,eg,eeg,ev,eev,egv,eegv,w,ww,wg,wwg,wv,wwv,wgv,wwgv Listpartsofyourcontactlist.Thefollowingflagsaredisplayedinthefirstcolumn: +Thisentryisactuallynorealcontact,butanaliasoftheprecedingone.Only withww. #Thisentryisnotonyourcontactlist,buttheUINwasusedatsometime.Only withwandww. *Thiscontactwillseeyourstatusevenifyou´ reinvisible. −Thiscontactwon´ tseeyouatall. ˆ Thiscontactisignored:nomessages,nostatuschanges. Theww*andee*groupofcommandsdisplayanothercolumnfordirectconnectionflags. Ifthefirstcolumnwouldbeempty,itwillbedisplayedinthefirstcontactforallofthose commandsunlessit´ sˆ . &Adirectconnectiontothiscontactisestablished. |Adirectconnectiontothiscontacthasfailed. :Adirectconnectioniscurrentlytriedtobeestablished. ˆ Nodirectconnectionopened,butIPaddressandportnumberareknown. Anothercolumnisaddedfortheww*andee*groupofcommandstoshowthestatus withregardstotheserversidecontactlist,asfarasitisknown: SThecontactisintheserversidecontactlist,andshouldbethere. sThecontactisintheserversidecontactlist,butshouldnotbethere. .Thecontactisnotintheserversidecontactlist,butshouldbethere.Thisusuallymeans"authorizationrequired". ´´ Thecontactisnotintheserversidecontactlist,andisnotwantedthere. climm 3 CLIMM(7) CLIMM(7) e[<group>] Listallpeoplewhichareonlineinyourcontactlist.Alsoprintyourownstatus. ee[<group>] Listallpeoplewhichareonlineinyourcontactlistwithmoredetails. eg,eeg Asabove,butsortbygroups.Newfor0.4.10. w,ww,wg,wwg Asabove,butalsoincludeofflinecontacts. ev,eev,egv,eegv,wv,wwv,wgv,wwgv Asabove,butalsoincludehiddencontacts(withtheshadowoptioneffectivelyset).New for0.5. ewideListallpeoplewhichareonlineinyourcontactlistinascreenwideformat. wideListallpeopleinyourcontactlistinascreenwideformat. search search<em@il> search<contact> search<first><last> Searchforauserwithem@ilastheiremailaddress,withnickastheirnick(which mustn´ tcontainan@),orwithfirstastheirfirstandlastastheirlastname.Ifnoargumentisgiven,askfornickname,firstname,lastname,emailaddressandawholebunch ofotherdatatosearchfor. rand[<number>] Findarandomuserininterestgroupnumber.Ifnumberisnotgiven,listallinterest groups. COMMANDSCONCERNINGYOURCONTACTLIST Note:youneedtosavetomakeanyofthechangesdonebythesecommmandspersistentforthe localcontactlist. add<uin><nickname> Adduintoyourcontactlistasnickname. add<contact><alias> addalias<contact><alias> Addaliasaliasfornickname.New(addalias)for0.4.10. add[<group>[<contacts>] addgroup<group>[<contacts>] Addallcontactsincontactstothecontactgroupgroup,whichiscreatedifitdoesn´ t existandthecommandisgivenasaddgroup.Newfor0.4.10. rem<contacts> Removethealiasesincontacts.Ifanynickisthelastaliasforthiscontact,removethe contactcompletely. Note:Thiscommandhasadifferentmeaningifthefirstargumentofthecontactsisa groupname,soifyouwanttoaffectallcontactsofagroup,youmaynotusethisasthe firstparameter;e.g.mentiononememberofthegroupfirst. remalias<contacts> Removethealiasesincontacts. Modifiedfor0.6.3. climm 4 CLIMM(7) CLIMM(7) remall<contacts> remcont<contacts> Removethecontactcompletelyfromthecontactlist. New(remcont)for0.6.3. Note:Thecommandremhasadifferentmeaningifthefirstargumentofthecontactsisa groupname,soifyouwanttoaffectallcontactsofagroup,youmaynotusethisasthe firstparameter;e.g.mentiononememberofthegroupfirst. rem<group><contacts> remgroup<group><contacts> Removeallcontactsincontactsfromthecontactgroupgroup.Modifiedfor0.6.3. remall<group> remgroupall<group> Removeallcontactsfromthecontactgroupgroup.Furtherargumentsmaybesilently ignored.Ifremgroupallisused,discardofthecontactgroupcompletely.Modifiedfor 0.6.3. togig<contacts> Toggleswhethercontact´ smessagesandstatuschangesareignored. toginv<contacts> Toggleswhetheryouwillbehiddenforcontacts. togvis<contacts> Toggleswhethercontactscanseeyouevenifyouareinvisible. COMMANDSCONCERNINGYOURICQACCOUNT pass<password> Changesyouricqpasswordtopassword. Note:yourpasswordmaynotstartwitho´ (thebyte0xf3). Note:youneedtosavetomakethispersistentincaseyousaveyourpasswordinyour ˜ /.climm/climmrcfile(seeclimmrc(5)),oryouwillgetamismatchedpasswordonyour nextlogin. update Updatesyourbasicuserinformation(email,nickname,etc.). otherUpdatesotheruserinformationlikeageandsex. aboutUpdatesyouraboutuserinformation. setr[<number>] Setsyourrandomusergrouptonumber.Withoutargument,listspossibleinterest groups. reg<password> Createsanewuseraccountwithpasswordpassword. COMMANDSFORCONFIGURINGTHECLIMMCLIENT verbose[<level>] Setverbosityofclimmtolevel,orshowthecurrentverbosity.Theverbositylevelisaset offlags: 8Showprotocolerrors. 16Showcreationanddeletionofpackets. 32Showaddedandremovedqueueevents. 64Showcreatedanddeletedconnections. 256Showv6packetsandtheirdata. climm 5 CLIMM(7) CLIMM(7) 4096Showv7/v8packets. 8192Showthehexdumpofv7/v8packets. 16284Savethosepacketstodisc. 65536Showpeer−to−peerpackets. 131072 Showthehexdumpofpeer−to−peerpackets. 262144 Savethosepacketstodisc. 2097152 Showpeer−to−peerhandshake. 4194304 ShowI/Oconnectionsetup. However,anynon−zeroverboselevelwillcausemoreinformationtobeprinted,andany levelgreaterthan1evenmore. clearClearthescreen. sound[on|off|event] Togglesoundonoroff,orcalltheeventscriptforbeeps.Newsyntaxfor0.4.10. prompt<userprompt> Setuserpromptinuserprompt.Thefollowingvariableswillbeparsedandmaybeused inpromptstring: %Ubyyourcurrentusername %nbyyourcurrentnick %Sbydescriptionofyourstatus %sbyshortdescriptionofyourstatus %Pbyservername %pbytypeofserver(icq8,jabber,msn...) %abylastsenduin/nick %rbylastreciveduin/nick %tbytimeinformatHH:MM:SS %Tbytimeinformatgiveninoptionpromptstrftime. %%% %[0−9]c byforegroundcolor %[0−9]C bybackgroundcolor %[0−1]b bold.on=1(default),off=0 %[0−1]u underline.on=1(default),off=0 %[0−1]i inverse;on=1(default),off=0.Reversesforegroundandbackgroundcolors. %dsetdefaultallcolorsandstyles climm 6 CLIMM(7) CLIMM(7) \bbackspace \rcarriagereturn \nnewline \thorizontaltab \eESC \\backslash Example: prompt %4c%b%p%d://%1c%b%n%3c/%2c%b%s%8c %t%7c%b>%6c%r%7c%b<%6c%b%a%7c%b> autoaway[<timeout>|on|off] Setthetimeouttotimeout,orswitchautoawayonusingthelastusedordefaultauto awaytime,orswitchautoawayoffsavingthecurrentusedautoawaytime.Atimeoutof 0isequivalenttooff.Ifnoargumentisgiven,thecurrentsettingisdisplayed. alias[auto[expand]][<alias>[<expansion>]] Makeanewaliasnamedaliaswhichissubstitutedwithexpansion.Ifnoaliasisgiven, listallcurrentaliases.Ifnoexpansionisgiven,listthecurrentalias.Otherwise,createa newalias.Iftheautoorautoexpandkeywordisgiven,thenewaliasisauto-expanding, thatis,itisimmediatelyexpandedwhenthespaceorreturnkeyispresseddirectlyafter enteringthealias. Ifthestring%sispresentinexpansion,eachoccurenceisreplacedwiththegivenargumentswhenthealiasisinvoked,otherwisetheywillbeappended.Ifthestring%ris presentinexpansion,eachoccurenceisreplacedbythenicknameor,ifthenickisnot available,theUIN,ofthecontactthatsentthelastmessagetoyou,ortheemptystringif thereisnone.Ifthestring%aispresentinexpansion,eachoccurenceisreplacedbythe nicknameor,ifthenickisnotavailable,theUIN,ofthecontactthatyousentthelast messageto,ortheemptystringifthereisnone. Newfor0.4.10.Auto-expandingaliasesnewfor0.5.0.4. unalias<alias> Deletethealiasnamedalias.Newfor0.4.10. lang[<lang>|<nr>]... Switchtothelanguageandencodinglang,givenbyisocode,oroutputtranslationstring numbernr.ToselectthedefaultlanguagegivenbytheenvironmentvariablesLANG, LCALLandLCMESSAGES,use!,autoordefault.Toselectnotranslationatall, use.,noneorunload.Togetdebugginginformation,prefixitwithdebug. Thefilesfrom˜ /.climm/i18n/*.i18ntakeprecedenceovertheglobalones,thefilesfrom BASE/i18n/*.i18btakeprecedenceoverthosefrom˜ /.climm/i18n/*.i18n,whilethe "funny"textstakeprecedenceoverthestandardones. Youmayspecifyasmanyargumentsasyouwish,e.g.toprintstringnumber117inGermanandthenswitchbacktodefaultlanguage,use"transde117default". uptime Showthetimeclimmhasbeenrunning,plussomestatistics. set<option><value> Setoptionoptiontoeitheronoroff,oranoptionspecificsetofvalues.optioncanbe: color:usecolors, delbs:interpretdeletecharactersasbackspace, climm 7 CLIMM(7) CLIMM(7) funny: usefunnymessages, auto:sendauto−replieswhenyou´ ree.g.away, prompt: Typeoftheprompt: useruseuserprompt.Seecommandpromptinclimmcmds(7),andoptions promptandpromptstrftimeinclimmrc(5), uinhavethelastnickintheprompt, simple simpleprompt,bydefaultis"climm>". autosave: automaticallysavetheclimmrc, autofinger: automaticallyfingernewUINs, linebreak: setthestyleforline−breakingmessagestosimple,tosimplyprintthemessage afterthenickandwrap,break,toaddalinebreakbeforeprintingeachincoming message,indent,toindentalllinesofamessagetothelevelafterthenick,or smart,toaddalinebreakonlyifthemessagewouldn´ tfitontheline. Newoptionsfor0.4.10.Optionshermit,log,logonoff,silent,tabsremovedfor0.5;usethe optcommandforthem. opt[<contact>|<contactgroup>|connection|global[<option>[<value>]]] Setoptionoptionforcontactcontact,contactgroupcontactgroup,forthecurrentconnectionorgloballytothevaluevalue,ordisplaythecurrentvalue,ordisplayalloptions andtheirsettings. Whenresolvingoptionsforcontacts,theoptionforthecontactitselfwillbeused,ifset; otherwise,ifthecontactbelongstoagroup,theoptionforthatgroup,ifset;otherwise,if thecontactbelongstoaserverconnection(i.e.isnotyetremoved),theoptionforthe serverconnection;otherwisetheglobaloption,ifset.Otherwise,0forintegeroptions, FALSEforbooleanoptions,thecolorstringfor"nocolor"unlessthiscolorstringis requested,ortheemptystringforstringoptions. Whenresolvingoptionsforcontactgroups,thesameisdoneexceptforthefirststep; optionsforconnectionsaretakenfromtheconnection,thenglobally;andglobaloptions, ofcourse,onlyglobally. Thefollowingoptionscurrentlyexist,withtypeandapplicability: colorschemeintegerglobal Thenumberofthecolorschemetouse.Setto0todisableandusethecolors definedbelow. colornonecolorglobal Thecolorstringtousefornormaltext.Thekeywordsblack,red,green,yellow,blue,magenta,cyan,white,none,andboldareunderstoodandreplaced bytheirANSIcontrolsequences.Notethatanycoloralreadyincludesnone,so specifyboldalwaysafterthecolor.Anyothertextisinterpretedascontrolcharacterstooutputverbatim. colorservercolorglobal Thecolorstringtouseforserverstuff. colorclientcolorglobal Thecolorstringtouseforclientstuff. climm 8 CLIMM(7) CLIMM(7) colorinvcharcolorglobal Thecolorstringtousewhendisplayingbyte−sequencesinvalidforthegiven encoding. colorerrorcolorglobal Thecolorstringtouseforerrors. colordebugcolorglobal Thecolorstringtousefordebugoutput. colorquotecolorglobal Thecolorstringtouseforquotedtext. webawarebooleanconnection Whetherthecurrentstatusshouldbevisibleontheweb. hideipbooleanconnection Whethertohidethe(local)IPaddress. dcauthbooleanconnection Whetherauthorizedcontactscanseethe(remote)IPaddress. dccontbooleanconnection Whethercontactscanseethe(remote)IPaddress. countawaybooleanconnection Ifset,incomingmessageswillbecountedandtheiroriginatorsshownifthestatus ismanuallychangedtoanythingotherthanonlineorffc(withorwithoutinvisible).Otherwise,thisisonlydoneifthestatuswasautomaticallychangedto awayorna(withorwithoutinvisible). obeysblbooleanconnection Whethertoloadtheserver−basedcontactlistuponlogin.Thiswillenforce requiredauthorizations,whichwillbeignoredforlocalcontactlists. ignorebooleancontact Whethertoignoreallmessagesfromthiscontact. hidefrombooleancontact Whethertobeinvisibleforthiscontact. intimatebooleancontact Whethertobeallwaysvisibleforthiscontact. logonoffbooleancontact Whethertologwhenthiscontactsgoesonlineoroffline. logchangebooleancontact Whethertologstatuschangesforthiscontact. logmessbooleancontact Whethertologmessagesforthiscontact. showonoffbooleancontact Whethertoshowwhenthiscontactgoesonlineoroffline. showchangebooleancontact Whethertoshowstatuschangesforthiscontact. autoautobooleancontact Whethertoautomaticallyrequestautomaticmessagesonstatuschangesforthis contact. hideackbooleancontact Whethertohideacknowledgesformessagestothiscontact. Note:usingthisisareallybadidea,asyou’llmisswhenmessagesareresent climm 9 CLIMM(7) CLIMM(7) withoutacknowledge. wantsblbooleancontact Whetheryouwantthiscontacttobeontheserversidecontactlist. peekmebooleancontact Whetheryouwantthiscontacttobepeekedwhenusingthepeekallorpeekall command. shadowbooleancontact Whetheryouwanttohidethiscontactincontactlistdisplaysaslongasnoneof the*vcommandsisused. localbooleancontact Whetherthiscontactisnotarealcontact,butjustanickforanumber. encodingstringcontact Theencodingtoassumeformessageswhoseencodingisundefined,inparticular metadataontheserver. tabspoolintegercontact Whethertoaddacontacttothetabslistuponstartup. autoawaystringcontact Theautoawaystringforstatusawaytoreturntothiscontact. autonastringcontact Theautoawaystringforstatusnatoreturntothiscontact. autooccstringcontact Theautoawaystringforstatusocctoreturntothiscontact. autodndstringcontact Theautoawaystringforstatusdndtoreturntothiscontact. autoffcstringcontact Theautoawaystringforstatusffctoreturntothiscontact. colormessagestringcontact Thecolorstringtouseformessagesofthiscontact. colorsentstringcontact Thecolorstringtouseforthenickwhenmessagesaresent. colorackstringcontact Thecolorstringtouseforthenickwhenmessagesareacknowledged. colorincomingstringcontact Thecolorstringtouseforthenickwhenmessagesarereceived. colorcontactstringcontact Thecolorstringtouseforthenickotherwise. Newfor0.5. optcontact<contact>[<option>[<value>]] Setordisplayoptionoptionoralloptionsofcontactcontact.Newfor0.5. optgroup<group>[<option>[<value>]] Setordisplayoptionoptionoralloptionsofcontactgroupgroup.Newfor0.5. optconnection[<option>[<value>]] Setordisplayconnectionoptionoption.Newfor0.5. optglobal[<option>[<value>]] Setordisplayglobaloptionoption.Newfor0.5. climm 10 CLIMM(7) CLIMM(7) saveSavescurrentsettingsintheconfigurationfile,whichis˜ /.climm/climmrc,unlessspecifiedotherwise.Beware,itwillclobberanycommentsinthisfile. q[<msg>] quit[<msg>] exit[<msg>] Quitclimm.Ifgiven,sendthemessagemsgtoallcontactspreviouslymessagedthatare stillonlineandonthecontactlist. Note:climmwillnotwaitforanyacknowledges-ifthemessageisnotreceivedbythe contactforwhateverreason,itwillnotberesent. Aliasesquitandexitnewfor0.4.10. x[<msg>] Quitclimmwithoutsaving.Ifgiven,sendthemessagemsgtoallcontactspreviouslymessagedthatarestillonlineandonthecontactlist. Note:climmwillnotwaitforanyacknowledges-ifthemessageisnotreceivedbythe contactforwhateverreason,itwillnotberesent. Aliasesquitandexitnewfor0.4.10.Newfor0.5. COMMANDSCONCERNINGADVANCEDFEATURES meta[show|load|save|set|get|rget]<contacts> Handlemetadataofcontacts.Thefollowingsubcommandsareknown: showShowsthemetadataofallcontactsgiven. loadLoadsthemetadataforallcontactsgivenfromdiscandshowsthem. saveSavesthemetadataforallcontactsgiventodisc. setUploadsyourmetadatatotheserver. getGetsthemetadataforallcontactsfromtheserverandshowsthem. getrGetsthemetadataforthecontactthelastmessagewasreceivedfromfromthe serverandshowsit. Newfor0.4.10. file[...] Aliasforpeerfile.Newfor0.4.10. accept[...] Aliasforpeeraccept.Newfor0.4.10.4. peer<command><uin|nick> OperatecommandcommandonusergivenbyUINuinornicknamenick. openOpenapeertopeerconnectionoverTCPtotheuser. closeCloseandresetapeertopeerconnectiontotheuser. offSwitchofftryingtoestablishsuchaconnectionforsendingmessagesuntilitis explicitlyopenedorreset. file<file><description> Sendfilefilewithdescriptiondescription. files[<file><as>]...<description> Sendfilestotheuser.Theremaybearbitrarilymanypairsofaphysicalfilename fileandthenametobepresentedtothepeer,as.Ifasis´ /´ ,thefilename withoutthepathissent,andifitis´ .´ thesamenameissent. accept[<contact>][<id>] AcceptanincomingfilerequestfromcontactwiththegivenIDid.Ifthereis onlyonependingincomingfilerequest,thecontactandidargumentsmaybe climm 11 CLIMM(7) CLIMM(7) omitted,otherwiseitisundefinedwhichfilerequestisansweredwhenseveral match. deny[<contact>][<id>][<reason>] DenyanincomingfilerequestfromcontactwiththegivenIDidwiththereason reason. login[...] Aliasforconnlogin.Newfor0.4.10.4. conn[<command><nr>] Listallconnections,oroperatecommandonconnectionnr. openOpenthegiven,orthefirstserverconnection.Anoptionalpasswordmaybe given,whichoverridesanypreviouslyknownone. loginOpenthegiven,orthefirstserverconnection.Anoptionalpasswordmaybe given,whichoverridesanypreviouslyknownone. closeClosethegivenconnection.Temporaryconnectionswillgetremovedbythis. remove Closeandremovegiven(temporary)connection. selectSelectthegivenserverconnectionasthecurrentone.nrmaybetheconnection numberortheUINusedfortheconnection. contact[<command>] Handletheserversidecontactlist: showDownloadtheserversidecontactlistandjustshowit. diffDownloadtheserversidecontactlistandshowonlycontacts(uin/nickpairs)that arenotinthelocalcontactlist. addDownloadtheserversidecontactlistandaddallcontactstothelocalone. upload Trytouploadlocalcontactstotheserversidecontactlist. download Downloadcontactsfromtheserversidecontactlist,butavoidmodifyingalready existingcontacts. import Downloadcontactsfromtheserversidecontactlist,modifyingexistingcontacts. Note:Thiswillalwaystrytouploadonlythosecontactswiththewantsblcontactoption.Doaoptglobalwantsblonifyouwanttouploadallofyourcontacts. peek[<contacts>][all][<contacts>] Checkforeachcontactwhetheritisactuallyonlineornot.Thespecialkeywordallwill checkuponallcontactswiththepeekmeoptioneffectivelyset. Note:ThisabusesabugintheICQprotocoltofigurethisoutandthuscanstopworking atanytime.Noadditionalinformationexceptonlineorofflinecanbefoundoutthisway. peek2<contacts> Builtinaliasforgetautoawaycontacts.Canbeusedtodetectpresenceofusersofsome clients,butthisitselfcanbedetectedbytheotheruser’sclient.climmsince0.5willnot bedetectedbythis,butwilldetecttriesfromotherclients. peekall[<contacts>] Builtinaliasforpeek<contacts>all. as<nr|uin><cmd> ExecuteclimmcommandcmdasiftheconnectionnumbernrorfortheUINuinwould bethecurrentone. climm 12 CLIMM(7) CLIMM(7) COMMANDSFORSCRIPTING tclscript<file> Executethetclscriptfile,whichmaybegivenrelativetotheclimmbasedirectory. tcl<string> Executethetclcommandstring.Typetclclimmhelptogetalistofclimm−specifictcl commands. SEEALSO climm(1),climmrc(5) AUTHOR ThismanpagewasoriginallycreatedbyJamesMorrison<ja2morrison@student.math.uwaterloo.ca>forareferencetoallinteractivecommandsinclimm.Ithasbeenadaptedtocurrent usagebyRu¨ digerKuhlmann. climm 13